OpenStudy (anonymous):

I need help figuring out the inverse function of f(x)=2+10x I kind of have an idea of what to do but I'm not entirely sure

OpenStudy (anonymous):

what is your idea?

OpenStudy (anonymous):

solve for x, then switch the variables

OpenStudy (isaiah.feynman):

@tiffanyr96 There are steps for finding the inverse of a function in every math book, memorize them once and for all.

OpenStudy (anonymous):

or on google

OpenStudy (anonymous):



OpenStudy (anonymous):

I looked on google but it didn't help much, neither did my flvs lesson, I know you have to solve for x but im not sure on how you make the switch, like what are you replacing with y

OpenStudy (anonymous):

replace y by x

OpenStudy (anonymous):

and x by y

OpenStudy (isaiah.feynman):

The switch is this way, Where ever is an x put y there and where ever there is a y put x there!

OpenStudy (anonymous):

ok so I have f(x)=2+10x to start with then I switch it to y=2+10x

OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

then you solve for x by subtracting 10 from both sides?

OpenStudy (anonymous):

no solve for x, first by subtracting 2 from both sides, then dividing by 10 from both sides, this will isolate and solves for x

OpenStudy (anonymous):

ahhhh ok I get it now, my bad ok so then it would turn into y-2=10x

OpenStudy (anonymous):

correct, then isolate x by division

OpenStudy (anonymous):

ok I got y-2/10 now heres where I got confused on what comes next

OpenStudy (anonymous):

you can reduce 2/10, to 1/5.. and the y is also divided by 10.. now it would be \[x = \frac{ y }{ 10 }-\frac{ 1 }{ 5 }\] now switch the y and x

OpenStudy (anonymous):

oh wait.... would \[f ^{-1}(x) = \frac{ x-2 }{ 10 }\] also be the correct answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

awesome, thanks so much for ur help :)

OpenStudy (anonymous):

no prob :)